# Multi-Entry Generalized Search Trees for Indexing Trajectories (Systems Paper)

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Query optimizationccs: Information systems Spatial-temporal systems
@@ 1. Introduction
MGiST and MSP-GiST store a more detailed representation of the actual trajectory in the index. A trajectory is represented by a collection of smaller sub-trajectories that can each be indexed individually in MGiST and MSP-GiST. One extreme is to index each trajectory segment (pair of subsequent observations) individually. This method produces the best accuracy (in terms of false positives) but creates indices that are larger than the actual data, which is not desirable due to their large storage cost and long querying time. As a trade-off, one could split a trajectory into a set of smaller sub-trajectories, and index the bounding boxes of these sub-trajectories. Splitting the trajectory into a set of smaller bounding boxes produces a better approximation of the actual trajectory than representing it with a single bounding box, which improves index efficiency by reducing the number of false positives (See Figure 1). By limiting the number of bounding boxes, one can bound the total size of the index in case of storage constraints.
@@
The rest of this paper proceeds as follows. Sections 2 and 3 present related work, and overviews the GiST and SP-GiST frameworks. Section 4 presents the MGiST and MSP-GiST frameworks and describes how they can be realize instances of multi-entry indices. As a case study, Section 5 presents the implementation of multi-entry R-Trees, Quad-Trees, and KD-Trees for trajectories, together with their trajectory splitting algorithms. Section 6 presents the experimental evaluation of MGiST- and MSP-GiST-based indices using synthetic and real-world data sets. Finally, Section 7 discusses possible future work, and Section 8 concludes the paper.
@@ 2. Related Work
GiST: To simplify the implementation of new indices, Hellerstein et al. (Hellerstein et al., 1995) introduce GiST, a generalized search tree that is easily extensible for different data types and queries. GiST can instantiate balanced search trees, e.g., the B-Tree, R-Tree or M-Tree, using minimal implementation effort through the use of pluggable modules. Behind the scenes, GiST transparently handles the complex index internals, e.g., tree balancing, concurrency, and recovery (Kornacker et al., 1997).
@@
SP-GiST: With the emergence of new database applications, new indexing solutions, e.g., KD-Trees, Quad-Trees, Tries, and other unbalanced search trees have become more prominent. However, these indices could not be implemented in GiST, as they are not balanced trees. Aref et al. (Aref and Ilyas, 2001) introduce SP-GiST, a space-partitioning generalized index, that can instantiate space-partitioning unbalanced trees in a generalized framework. Many improvements and optimizations have been proposed for both GiST and SP-GiST, e.g.,  (Aoki, 1998; Ghanem et al., 2004; Beier et al., 2012) as well as their implementation in existing database systems, e.g., (Eltabakh et al., 2006; Borodin et al., 2017). GiST and SP-GiST are further detailed in Section 3.
@@
The Temporal Dimension: Many specialized indices have also been developed for the indexing of spatial and spatio-temporal data. The R-Tree (Guttman, 1984) indexes spatial objects using their bounding boxes in a balanced tree structure. By storing the time as an additional dimension, spatio-temporal objects can be indexed in 3D R-Trees (Theodoridis et al., 1996) or RT-Trees (Xu et al., 1990). Another approach to index the temporal dimension is to have an R-Tree for each time instant. To save space, consecutive R-Trees can also share sub-trees that remain unchanged between two time instants, e.g., as in (Xu et al., 1990; Nascimento and Silva, 1998; Tao and Papadias, 2001). There are also combinations of the above approaches, e.g., as in (Xu et al., 1990).
@@
Indexing Trajectories: A trajectory is a special type of spatio-temporal data. It combines multiple trajectory segments into a single large data object, with each segment being formed by linear interpolation of two stored instants. In the case of a primary index, each segment needs to be stored in the index, e.g., in a 3D R-Tree. STR-Trees and TB-Trees (Pfoser et al., 2000) are extensions of the 3D R-Tree that cluster segments of the same trajectory to improve trajectory-related queries. If the tree is used as a secondary index, not all the segments have to be stored individually. Thus, one can index fewer bounding boxes that each represents a subset of the complete trajectory. This is the scenario handled by this paper. Hadjieleftheriou et al. (Hadjieleftheriou et al., 2002) and Rasetic et al. (Rasetic et al., 2005) present possible splitting algorithms for this purpose. These algorithms are re-used in the implementation of MGiST and MSP-GiST for trajectories as in Section 5.
@@ 3.1. GiST Key Methods
PickSplit($P$): Given a set $P=\{E_{1},...,E_{M+1}\}$ of $M+1$ entries, splits $P$ into two subsets $P_{1}$ and $P_{2}$, each containing at least $k*M$ entries, where $M$ is the maximum node capacity of the tree and $k$ corresponds to the minimum fill factor.
@@ 3.2. SP-GiST Key Methods and Parameters
Similar to GiST, SP-GiST presents a list of key methods that are to be implemented to instantiate a specific space partitioning index. Methods Consistent and PickSplit are the same as those of GiST.
@@ 4.3. Nearest Neighbour Search
In this section, we discuss the similarities and differences between MGiST and MSP-GiST in one side and another generalized index structure: GIN (gin, 2024) on the other side. We further explain the use cases of each. GIN is a generalized inverted index that can be used to index arrays and documents (gin, 2024). Its implementation also requires an ExtractValue method that also splits a tuple into multiple sub-elements, termed keys (e.g., values in an array, or words in a document). This is the main similarity. However, two main structural differences between MGiST/MSP-GiST and GIN indices. (1) GIN indexes the keys in a B-Tree. This makes GIN not suitable for indexing complex spatial and spatio-temporal objects. (2) The data indexed by GIN is assumed to contain many duplicate keys. For example, a word can be present in many different documents. To handle this, GIN indexes the distinct keys in a B-Tree and stores for each key at the leaf level a set of tuple pointers that contain this key. In MGiST and MSP-GiST, entries with the same keys are stored as separate entries in the index leaves. Adding this de-duplication of keys/predicates to MGiST and MSP-GiST is left as future work.
@@ 5. Multi-Entry Trajectory Indexing
(5) $p(t)=\frac{p_{i}*(t_{i+1}-t)+p_{i+1}*(t-t_{i})}{t_{i+1}-t_{i}},t_{i}\leq t<t_{i+1}$
There are many ways to split a trajectory into multiple bounding boxes. These methods vary in the way the splits are generated and in how many boxes the trajectory is split. One straightforward splitting algorithm is the following. Given a trajectory $r$ and a constant parameter $k$ (the same for all tuples), split $r$ into $k$ ‘equal’ parts. This method produces a box for every $\lceil\frac{n}{k}\rceil$ consecutive segments. Refer to this splitting algorithm by EquiSplit. Notice that EquiSplit splits all trajectories into the same number of $k$ boxes. Given a dataset of $N$ trajectories, EquiSplit produces in total $N*k$ boxes, i.e., the size of the created index is directly proportional to $k$.
@@
In both algorithms, the number of bounding boxes per trajectory is constant. This is a big restriction and is not optimal. In fact, in case some trajectories are much longer than others (have a bigger spatial extent), it could be beneficial to split them into more boxes. Thus, we propose an algorithm, termed ManualSplit, that requires a parameter $m$; the number of segments that form a bounding box. A trajectory with $n$ segments is split into $\lceil\frac{n}{m}\rceil$ boxes. ManualSplit is the inverse of EquiSplit, which has a fixed number of boxes per trajectory but a variable number of segments per box.
@@
Rasetic et al. (Rasetic et al., 2005) present a cost model for splitting trajectories based on the expected number of I/O’s of a given range query. They present an optimal splitting algorithm that minimizes the cost of a given range query. To avoid quadratic time complexity, they propose a heuristic linear-time algorithm, termed LinearSplit, that produces near-optimal results. LinearSplit starts by linearly accumulating the trajectory instants until a given criterion is met. A constant-slope approximation of the collected sub-sequence is used to compute the bounding boxes covering this part of the trajectory. This step takes linear time. LinearSplit continues with the remaining trajectory instants. For more detail, the readers are referred to (Rasetic et al., 2005). Similarly to ManualSplit and AdaptSplit, LinearSplit produces a variable number of boxes per trajectory.
(a) EquiSplit / ManualSplit
@@
Figure 2 illustrates the result of applying the different splitting algorithms on a single trajectory. Each image contains two splits, one in 7 bounding boxes (with solid border) and one in 28 boxes (dotted border). Note that for a single trajectory, EquiSplit and ManualSplit are equivalent. Indeed, given a number of boxes $k$ and a trajectory $T$ of length $n$, one can compute the parameter $m=\lceil\frac{n}{k}\rceil$, such that EquiSplit($T$, $k$) and ManualSplit($T$, $m$) produce the same boxes. The same holds for MergeSplit and AdaptSplit. However, this is not true anymore when applying the algorithm with the same parameter on multiple trajectories of different lengths.
@@ 6.1. The BerlinMOD Benchmark
BerlinMOD (Düntgen et al., 2009) is a benchmark for moving objects databases that presents both a synthetic trajectory data generator and a set of benchmark queries. In these experiments, we use a version of the BerlinMOD data generator that is implemented in MobilityDB, and thus generates the trajectories readily in MobilityDB moving point format. [^1] https://github.com/MobilityDB/MobilityDB-BerlinMOD The generator simulates work and leisure trips in and around the city of Brussels. It is parameterized by a scale factor that determines the size of the generated data set. We use a scale factor of 1 that produces trips for 2000 vehicles over 30 days, for a total of 157K trips containing on average 1.2K instants each. The size of this trajectory data set is 7.7GB. Next to the trajectory data, the generator produces four other tables of interest: QueryInstants, QueryPeriods, QueryPoints and QueryRegions, that are used in the benchmark queries. Table 3 summarizes the generated tables.
@@
Trips (vehicle_id, trip_id, trip) 157549
QueryInstants (id, instant) 100
QueryPeriods (id, period) 100
QueryPoints (id, point) 100
QueryRegions (id, region) 100
Table 2. Tables produced by the BerlinMOD generator.
@@
A spatio-temporal index is constructed on the trip column of the Trips table, for different parameter values of the splitting algorithms. The parameter for MergeSplit is the number of boxes per trajectory. For ManualSplit and AdaptSplit, the parameter is the (average) number of segments per box. Lastly, the parameter for LinearSplit corresponds to the average width, length and duration, in meters and minutes respectively, of the query box.
Figure 4 gives the index size and construction time for the five splitting algorithms. The experiments confirm linear proportionality between the number of splits and the index size in the case of MergeSplit, as mentioned in Section 5, and negative exponential in the case of ManualSplit and AdaptSplit. For LinearSplit, a decrease in average query size results in an increase in number of generated boxes and thus an increase in index size.
@@ 6.1.1. Point and Range Queries
The least expensive query is Query 11. This query applies a point intersection test at a snapshot of the returned trajectories. For this query, the multi-entry R-Tree implementations is actually slower than its traditional single-entry counterparts. In this case, the index scan is dominating the query duration, and is thus slowing down the query. The same effect can be seen for the KD-Tree at large index sizes, but to a lesser extent. The multi-entry MSP-GiST indices thus seem to have a lower index scan cost, as they are still able to improve the duration of Query 11.
@@ 6.1.2. Nearest-Neighbor Search Queries
    ORDER BY trip |=| stbox(p.geom)
@@
These queries find the 5 trajectories with the closest point of approach to the query point (Query 1/NN) or the spatio-temporal points constructed by joining the query points and the query periods (Query 2/NN). The first query is essentially a spatial-only KNN query, while the second query is applied on temporal ranges.
@@ 6.1.3. Comparison with Single-Entry Indices
Concerning the splitting algorithms, MergeSplit is consistently performing the worst. This is due to the fact that it splits each trajectory into an equal number of bounding boxes. This is not optimal when working with trajectories of various sizes. Secondly, the three remaining algorithms are approximately equivalent on Queries 4 and 13, with ManualSplit performing slightly worse on Query 11. We further evaluate these three algorithms on real-world trajectory data in Section 6.2.
@@ 6.2. Vessel Trajectory Data
In this section, we further evaluate the performance of the multi-entry Quad-Tree on a real-world data set. The Danish Maritime Authority gives access to historical AIS data around Denmark. [^2] http://web.ais.dk/aisdata/ In this experiment, we use the AIS data for the month of January 2023. The raw data set is 54.4GB. After importing the data into MobilityDB and applying basic data cleaning, the remaining data set contains 53K trajectories having on average 2.9K instants each. We also manually construct two additional tables: a table containing 10 ports around Denmark, and a table containing the periods from 2pm to 4pm for Week 1 of January 2023. Table 3 lists these tables.
@@
Ships (mmsi, trip_id, trip) 53008
Ports (port, region) 10
Periods (day, period) 7
Table 3. Tables used for the AIS experiment.
We construct multi-entry Quad-Tree indices on the trip column of the Ships table, for the ManualSplit, AdaptSplit and LinearSplit algorithms. We evaluate the following 2 range queries:

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3678717.3691320conference: The 32nd ACM International Conference on Advances in Geographic Information Systems; October 29-November 1, 2024; Atlanta, GA, USAbooktitle: The 32nd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’24), October 29-November 1, 2024, Atlanta, GA, USAisbn: 979-8-4007-1107-7/24/10ccs: Information systems Query optimizationccs: Information systems Spatial-temporal systems
@@ 1. Introduction
MGiST and MSP-GiST store a more detailed representation of the actual trajectory in the index. A trajectory is represented by a collection of smaller sub-trajectories that can each be indexed individually in MGiST and MSP-GiST. One extreme is to index each trajectory segment (pair of subsequent observations) individually. This method produces the best accuracy (in terms of false positives) but creates indices that are larger than the actual data, which is not desirable due to their large storage cost and long querying time. As a trade-off, one could split a trajectory into a set of smaller sub-trajectories, and index the bounding boxes of these sub-trajectories. Splitting the trajectory into a set of smaller bounding boxes produces a better approximation of the actual trajectory than representing it with a single bounding box, which improves index efficiency by reducing the number of false positives (Fig. 1). By limiting the number of bounding boxes, one can bound the total size of the index in case of storage constraints.
@@
The rest of this paper proceeds as follows. Sections 2 and 3 present related work, and overviews the GiST and SP-GiST frameworks. Section 4 presents the MGiST and MSP-GiST frameworks and describes how they can realize instances of multi-entry indices. As a case study, Section 5 presents the implementation of multi-entry R-Trees, Quad-Trees, and KD-Trees for trajectories, together with their trajectory splitting algorithms. Section 6 presents the experimental evaluation of MGiST- and MSP-GiST-based indices using synthetic and real-world data sets. Finally, Section 7 discusses possible future work, and Section 8 concludes the paper.
@@ 2. Related Work
To simplify the implementation of new indices, Hellerstein et al. (Hellerstein et al., 1995) introduce GiST, a generalized search tree that is easily extensible for different data types and queries. GiST can instantiate balanced search trees, e.g., the B-Tree, R-Tree or M-Tree, using minimal implementation effort through the use of pluggable modules. Behind the scenes, GiST transparently handles the complex index internals, e.g., tree balancing, concurrency, and recovery (Kornacker et al., 1997).
SP-GiST
@@
With the emergence of new database applications, new indexing solutions, e.g., KD-Trees, Quad-Trees, Tries, and other unbalanced search trees have become more prominent. However, these indices could not be implemented in GiST, as they are not balanced trees. Aref et al. (Aref and Ilyas, 2001) introduce SP-GiST, a space-partitioning generalized index, that can instantiate space-partitioning unbalanced trees in a generalized framework. Many improvements and optimizations have been proposed for both GiST and SP-GiST, e.g.,  (Aoki, 1998; Ghanem et al., 2004; Beier et al., 2012) as well as their implementation in existing database systems, e.g., (Eltabakh et al., 2006; Borodin et al., 2017). GiST and SP-GiST are further detailed in Section 3.
The Temporal Dimension
@@
Many specialized indices have also been developed for the indexing of spatial and spatio-temporal data. The R-Tree (Guttman, 1984) indexes spatial objects using their bounding boxes in a balanced tree structure. By storing the time as an additional dimension, spatio-temporal objects can be indexed in 3D R-Trees (Theodoridis et al., 1996) or RT-Trees (Xu et al., 1990). Another approach to index the temporal dimension is to have an R-Tree for each time instant. To save space, consecutive R-Trees can also share sub-trees that remain unchanged between two time instants, e.g., as in (Xu et al., 1990; Nascimento and Silva, 1998; Tao and Papadias, 2001). There are also combinations of the above approaches, e.g., as in (Xu et al., 1990).
Indexing Trajectories
@@
A trajectory is a special type of spatio-temporal data. It combines multiple trajectory segments into a single large data object, with each segment being formed by linear interpolation of two stored instants. In the case of a primary index, each segment needs to be stored in the index, e.g., in a 3D R-Tree. STR-Trees and TB-Trees (Pfoser et al., 2000) are extensions of the 3D R-Tree that cluster segments of the same trajectory to improve trajectory-related queries. If the tree is used as a secondary index, not all the segments have to be stored individually. Thus, one can index fewer bounding boxes that each represents a subset of the complete trajectory. This is the scenario handled by this paper. Hadjieleftheriou et al. (Hadjieleftheriou et al., 2002) and Rasetic et al. (Rasetic et al., 2005) present possible splitting algorithms for this purpose. These algorithms are re-used in the implementation of MGiST and MSP-GiST for trajectories as in Section 5.
@@ 3.1. GiST Key Methods
PickSplit($P$): Given a set $P=\{E_{1},...,E_{M+1}\}$ of $M+1$ entries, splits $P$ into two subsets $P_{1}$ and $P_{2}$, each containing at least $k\cdot M$ entries, where $M$ is the maximum node capacity of the tree and $k$ corresponds to the minimum fill factor.
@@ 3.2. SP-GiST Key Methods and Parameters
Similar to GiST, SP-GiST presents a list of key methods that are to be implemented to instantiate a specific space partitioning index. Methods Consistent and PickSplit are similar to those of GiST.
@@ 4.3. Nearest Neighbour Search
In this section, we discuss the similarities and differences between MGiST and MSP-GiST in one side and another generalized index structure: GIN (gin, 2024) on the other side. We further explain the use cases of each. GIN is a generalized inverted index that can be used to index arrays and documents (gin, 2024). Its implementation also requires an ExtractValue method that also splits a tuple into multiple sub-elements, termed keys (e.g., values in an array, or words in a document). This is the main similarity. However, two main structural differences exist between MGiST/MSP-GiST and GIN indices. (1) GIN indexes the keys in a B-Tree. This makes GIN not suitable for indexing complex spatial and spatio-temporal objects. (2) The data indexed by GIN is assumed to contain many duplicate keys. For example, a word can be present in many different documents. To handle this, GIN indexes the distinct keys in a B-Tree and stores for each key at the leaf level a set of tuple pointers that contain this key. In MGiST and MSP-GiST, entries with the same keys are stored as separate entries in the index leaves. Adding this de-duplication of keys/predicates to MGiST and MSP-GiST is left as future work.
@@ 5. Multi-Entry Trajectory Indexing
(5) $p(t)=\frac{p_{i}\cdot(t_{i+1}-t)+p_{i+1}\cdot(t-t_{i})}{t_{i+1}-t_{i}},t_{i}\leq t<t_{i+1}$
There are many ways to split a trajectory into multiple bounding boxes. These methods vary in the way the splits are generated and in how many boxes the trajectory is split. One straightforward splitting algorithm is the following. Given a trajectory $r$ and a constant parameter $k$ (the same for all tuples), split $r$ into $k$ ‘equal’ parts. This method produces a box for every $\lceil\frac{n}{k}\rceil$ consecutive segments. Refer to this splitting algorithm by EquiSplit. Notice that EquiSplit splits all trajectories into the same number of $k$ boxes. Given a dataset of $N$ trajectories, EquiSplit produces in total $N\cdot k$ boxes, i.e., the size of the created index is directly proportional to $k$.
@@
In both algorithms, the number of bounding boxes per trajectory is constant. This is a big restriction and is not optimal. In fact, in case some trajectories are much longer than others (have a bigger spatial extent), it could be beneficial to split them into more boxes. Thus, we propose an algorithm, termed SegSplit, that requires a parameter $m$; the number of segments that form a bounding box. A trajectory with $n$ segments is split into $\lceil\frac{n}{m}\rceil$ boxes. SegSplit is the inverse of EquiSplit, which has a fixed number of boxes per trajectory but a variable number of segments per box.
@@
Rasetic et al. (Rasetic et al., 2005) present a cost model for splitting trajectories based on the expected number of I/O’s of a given range query. They present an optimal splitting algorithm that minimizes the cost of a given range query. To avoid quadratic time complexity, they propose a heuristic linear-time algorithm, termed LinearSplit, that produces near-optimal results. LinearSplit starts by linearly accumulating the trajectory instants until a given criterion is met. A constant-slope approximation of the collected sub-sequence is used to compute the bounding boxes covering this part of the trajectory. This step takes linear time. LinearSplit continues with the remaining trajectory instants. For more detail, the readers are referred to (Rasetic et al., 2005). Similarly to SegSplit and AdaptSplit, LinearSplit produces a variable number of boxes per trajectory.
(a) EquiSplit / SegSplit
@@
Figure 2 illustrates the result of applying the different splitting algorithms on a single trajectory. Although the figure only illustrates the spatial projection of the bounding boxes, it’s important to keep in mind that all splitting algorithms also consider the temporal dimension. Each image contains two splits, one in 7 bounding boxes (with solid border) and one in 28 boxes (dotted border). Note that for a single trajectory, EquiSplit and SegSplit are equivalent. Indeed, given a number of boxes $k$ and a trajectory $T$ of length $n$, one can compute the parameter $m=\lceil\frac{n}{k}\rceil$, such that EquiSplit($T$, $k$) and SegSplit($T$, $m$) produce the same boxes. The same holds for MergeSplit and AdaptSplit. However, this is not true anymore when applying the algorithm with the same parameter on multiple trajectories of different lengths.
@@ 6.1. The BerlinMOD Benchmark
BerlinMOD (Düntgen et al., 2009) is a benchmark for moving objects databases that presents both a synthetic trajectory data generator and a set of benchmark queries. In these experiments, we use a version of the BerlinMOD data generator that is implemented in MobilityDB, and thus generates the trajectories readily in MobilityDB moving point format. [^1] https://github.com/MobilityDB/MobilityDB-BerlinMOD The generator simulates work and leisure trips in and around the city of Brussels. It is parameterized by a scale factor that determines the size of the generated data set. We use a scale factor of 1 that produces trips for 2000 vehicles over 30 days, for a total of 157K trips containing on average 1.2K instants each. The size of this trajectory data set is 7.7GB. Next to the trajectory data, the generator produces four other tables of interest: QueryInstants, QueryPeriods, QueryPoints and QueryRegions, that are used in the benchmark queries. The query regions (Fig. 3(a), black) have an average size of 1 km2. Table 2 summarizes the generated tables.
Table 2. Tables produced by the BerlinMOD generator.
@@
[…]
@@
A spatio-temporal index is constructed on the trip column of the Trips table, for different parameter values of the splitting algorithms. The parameter for MergeSplit is the number of boxes per trajectory. For SegSplit and AdaptSplit, the parameter is the (average) number of segments per box. Lastly, the parameter for LinearSplit corresponds to the average width, length and duration, in meters and minutes respectively, of the query box.
Figure 4 gives the index size and construction time for the four splitting algorithms. The experiments confirm linear proportionality between the number of splits and the index size in the case of MergeSplit, as mentioned in Section 5, and negative exponential in the case of SegSplit and AdaptSplit. For LinearSplit, a decrease in average query size results in an increase in number of generated boxes and thus an increase in index size.
@@ 6.1.1. Point and Range Queries
The least expensive query is Query 11. This query applies a point intersection test at a snapshot of the returned trajectories. For this query, the multi-entry R-Tree implementations is actually slower than its traditional single-entry counterparts. In this case, the index scan is dominating the query duration, and is thus slowing down the query. The same effect can be seen for the KD-Tree at large index sizes, but to a lesser extent. The multi-entry MSP-GiST indices thus seem to have a lower index scan cost, as they are still able to improve the duration of Query 11. The increased index scan cost of the R-Tree indices is potentially due to the sequential insert method used by MGiST producing R-Trees with high overlap between nodes.
@@ 6.1.2. Nearest-Neighbor Search Queries
    ORDER BY trip |=| stbox(point)
@@
These queries find the 5 trajectories with the closest point of approach to the query point (Query 1/NN) or the spatio-temporal points constructed by joining the query points and the query periods (Query 2/NN). The first query is essentially a spatial-only KNN query, while the second query is applied on temporal ranges. For clarity, we set $k$ to 5, though tests with $k$ up to two orders of magnitude larger yielded similar performance gains.
@@ 6.1.3. Comparison with Single-Entry Indices
Concerning the splitting algorithms, MergeSplit is consistently performing the worst. This is due to the fact that it splits each trajectory into an equal number of bounding boxes. This is not optimal when working with trajectories of various sizes. Secondly, the three remaining algorithms are approximately equivalent on Queries 4 and 13, with SegSplit performing slightly worse on Query 11. We further evaluate these three algorithms on real-world trajectory data in Section 6.2.
@@ 6.2. Vessel Trajectory Data
In this section, we further evaluate the performance of the multi-entry Quad-Tree on a real-world data set. The Danish Maritime Authority gives access to historical AIS data around Denmark. [^2] http://web.ais.dk/aisdata/ In this experiment, we use the AIS data for the month of January 2023. The raw data set is 54.4GB. After importing the data into MobilityDB and applying basic data cleaning, the remaining data set contains 53K trajectories having on average 2.9K instants each. We also manually construct two additional tables: a table containing 10 ports around Denmark (Fig. 3(b), black) with and average size of 11.5 km2, and a table containing the periods from 2pm to 4pm for Week 1 of January 2023. Table 3 lists these tables.
Table 3. Tables used for the AIS experiment.
@@
[…]
We construct multi-entry Quad-Tree indices on the trip column of the Ships table, for the SegSplit, AdaptSplit and LinearSplit algorithms. We evaluate the following 2 range queries: